[Hallc_running] [New Logentry] Hall C RC daily update

simona at jlab.org simona at jlab.org
Fri Jul 31 10:20:07 EDT 2020


Logentry Text:
--
Hall C was swept yesterday around 3:30 pm and then put in controlled access.

The BSY3C power supply could not be fixed yesterday :https://logbooks.jlab.org/entry/3818384 
The decision was made at the 8 am meeting this morning to replace this power supply with a spare, new one. If everything goes well we would be on track for tune beam to Hall C tomorrow during day shift.

Meanwhile Ops will set up Hall B and D for beam.

Hall C was given beam authorization up to 45 muA CW: https://logbooks.jlab.org/entry/3818564

Today's day, swing and tomorrow's owl shifts are cancelled: https://misportal.jlab.org/mis/physics/shiftSchedule/?experimentRunId=HALLC-POL3HE

I took an access this morning to check on the downstream beamline turbo pump and it's still running and the vacuum looks good: https://logbooks.jlab.org/entry/3818583

We will meet today at 4:15 pm. Here is how to connect via bluejeans:  https://bluejeans.com/861439920

---

This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3818585
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200731/86ecfd81/attachment.html>


More information about the Hallc_running mailing list